# Near-Optimal Channel Estimation for Dense Array Systems

### v1

@@ Abstract
By deploying a large number of antennas with sub-half-wavelength spacing in a compact space, dense array systems (DASs) can fully unleash the multiplexing-and-diversity gains of limited apertures. To acquire these gains, accurate channel state information acquisition is necessary but challenging due to the large antenna numbers. To overcome this obstacle, this paper reveals that exploiting the high spatial correlation of DAS channels is crucial while designing the observation matrix for optimal/near-optimal channel estimation. Firstly, we prove that the observation matrix design is equivalent to a time-domain duality of multiple-input multiple-output precoding, which can be ideally addressed by the water-filling principle. For practical realizations, a novel ice-filling algorithm is proposed to design amplitude-and-phase controllable observation matrices, and a majorization-minimization algorithm is proposed to address the phase-only controllable case. Particularly, we prove that the ice-filling algorithm can be viewed as a “quantized” water-filling algorithm. To support the sub-optimality of the proposed designs, we provide comprehensive analyses on the achievable mean square errors and their asymptotic expressions. Finally, numerical simulations verify that our proposed channel estimation designs can achieve the near-optimal performance and outperform existing approaches significantly.
@@ I Introduction
From 3G to 5G, antenna array systems play an irreplaceable role in wireless communications [1, 2, 3]. By strategically configuring multiple antennas, an antenna array can constructively manipulate the radiated signals, allowing for enhanced wireless coverage, improved spectral efficiency, and reduced power consumption [4, 5, 6]. The performance of arrays continuously improves with the number of antennas [7, 8]. However, in practical scenarios, the space available for antenna deployment is usually limited, which restricts the performance gains endowed by antenna arrays [9]. To achieve a better performance with a spatially-limited aperture, dense array systems have attracted extensive attentions in recent years [10].
@@
Generally, DASs represent a series of array technologies that massive sub-wavelength antennas are densely arranged within a compact space. Unlike the conventional arrays with half-wavelength antenna spacing $\lambda/2$, the antenna spacing of DASs is much smaller, such as $\lambda/6$ [11], $\lambda/8$ [12], $\lambda/10$ [13], or even $\lambda/23$ [14]. With an increased number of sub-channels, DASs promise to achieve high array gains and fully exploit the multiplexing-and-diversity gains of limited apertures [10]. Besides, DASs can also reduce the grating lobes and provide high performance for large values of oblique angles of incidence [15]. Some works also report their advantages of super-bandwidth and super-directivity [16, 17, 18]. The typical DAS realizations include holographic multiple-input multiple-output (MIMO) [19], reconfigurable intelligent surfaces [20, 21], reconfigurable holographic surfaces [22], fluid antenna systems [23, 24], graphene-based nano-antenna arrays [25], and so on. For example, in [11], massive sub-wavelength patches of $\lambda/6$-spacing are densely printed on a holographic MIMO surface to generate multiple beams flexibly. In [26], meta-elements sized of $\lambda/2.5\times\lambda/17.5$ are seamlessly integrated onto a feeding microstrip to achieve RHS beam steering. In [14], an RIS composed of antennas sized of $\lambda/23\times\lambda/23$ is designed and fabricated for terahertz communications.
@@
The performance gains of antenna arrays are realized via the constructive beamformers enabled by their phase shifters and radio frequency (RF) chains. To realize effective beamforming and signal demodulation, the acquisition of accurate channel state information (CSI) is essential. Up to now, many estimators have been proposed to acquire the CSI of large antenna arrays. Their fundamental principle involves receiving pilot signals using observation matrices and recovering the channel by advanced estimation algorithms. For example, when the available pilot length is larger than the antenna number, the classical non-parametric algorithms are widely used, including least square (LS) estimator and minimum mean square error (MMSE) estimator. Leveraging the property of channel sparsity, compressed sensing (CS)-based channel estimators are widely studied to improve the estimation accuracy and reduce the pilot overhead [27, 28], such as the orthogonal matching pursuit (OMP)-based estimator [29], the message passing (MP)-based estimator [30, 31, 32], and the gridless sparse signal reconstructor [33]. By training neural networks with a large amounts of channel data, the deep learning approaches are also utilized to realize data-driven channel estimators [34, 35]. Besides, beam alignment techniques [36, 37, 38], including beam sweeping and hierarchical beam training, have also been widely explored to acquire the implicit CSI with low pilot overhead.
@@
Although most of existing channel estimators can be adopted in DASs, they fail to fully exploit the high spatial correlation of DAS channels, thereby leaving a remarkable performance gap from the optimal estimator. Specifically, this spatial correlation is attributed to the fact that, the extremely-dense deployment of DAS antennas significantly increases the similarity of radio waves impinging on antenna ports and aggravates the mutual-coupling effect between adjacent antenna circuits. From the mathematical perspective, this fact implies that the covariance matrices of DAS channels are no longer diagonal, but highly structured and/or under-determined. For diagonal covariance matrices, a uniform or random pilot-sensing strategy is sufficient to achieve the optimality of estimation [39, 40]. Thereby, in conventional channel estimations, the observation matrices for receiving the pilot signals are either generated randomly or set as predefined codebooks, such as the Discrete Fourier Transform (DFT) matrix and the identity matrix [27, 28, 29, 33, 30, 31, 34, 35, 36, 37]. In contrast, since the correlation matrix of DAS channels are full of structural features, it is believed that their observation matrices can be tightly aligned with these features in pilot transmission. This structured pilot-sensing process has a high potential to remarkably boost the channel estimation accuracy in DASs [41, 42, 43].
@@
To achieve this goal, our work represents the first attempt on designing and analyzing the optimal/near-optimal observation matrices in the DAS channel estimation. By incorporating a mutual-information-maximization (MIM)-enabled observation matrix design into a Bayesian regression-based estimator, we propose a novel framework for DAS channel estimation. Our main contributions are summarized as follows.
@@
  • Framework of DAS channel estimator: From the perspective of MIM, the proposed framework of DAS channel estimation establishes the connection between channel estimation and MIMO precoding. Specifically, we propose to design the DAS’s observation matrix by minimizing the information uncertainty between the received pilots and wireless channels. The formulated MIM problem for observation matrix design is shown to be a time-domain duality of MIMO precoding design. In particular, the multi-RF-chain resources for precoding design is transformed to the multi-timeslot-pilot resources for channel estimation. Motivated by this finding, we prove that the ideal (but maybe unachievable in practice) observation matrix follows the water-filling principle, which is built by the eigenvectors of the channel’s covariance matrix followed by the optimal power allocation.
@@
  • Majorization minimization (MM) enabled observation matrix design: We then apply our framework for DAS channel estimation to analog combining/beamforming architecture. In this context, only the phases of the observation matrix’s weights are controllable while their amplitudes are fixed, making both the ice-filling and water-filling algorithms invalid. To address this challenge, we propose a MM algorithm for designing the observation matrix. Its novelty lies in replacing the primal non-convex MIM problem with a series of tractable approximate subproblems having analytical solutions, which are solved in an alternating optimization manner. Numerical simulations demonstrate that its channel estimation performance is comparable to the sub-optimal ice-filling algorithm.
  • Performance analysis: Comprehensive analyses on the achievable mean square errors (MSEs) are provided to validate the effectiveness of the proposed designs. We analytically prove that, compared to the random observation matrix design, the ice-filling algorithm can significantly improve the estimation accuracy by the order of the ratio between the number of antennas and the rank of channel covariance. Furthermore, based on the derived quantization nature of ice-filling, the MSE gap between water-filling and ice-filling is proved to decay quadratically with the pilot length, demonstrating the near-optimality of the ice-filling algorithm. In addition, the close-form MSEs under imperfect channel covariance are also derived. The superiority of the proposed algorithms is further validated using numerical results under both perfect and imperfect channel covariance conditions.
The remainder of the paper is organized as follows. The system model and problem formulation are presented in Section II. The proposed framework of channel estimation design is given in Section III. The practical observation matrix designs are elaborated in Section IV. Then, the MSEs are analyzed in Section V. Simulation results are carried out in Section VI, and finally the conclusions are drawn in Section VII.
@@ II System Model
Figure 1: An illustration of channel estimation for an DAS, where one $M$-antenna BS receives the pilots from a user in the uplink.
This paper commences by elaborating on the system model. As illustrated in Fig. 1, we focus on the narrowband channel estimation of an uplink single-input multiple-output (SIMO) system, which comprises an $M$-antenna BS equipped with one RF chain and a single-antenna user [7].
Let ${\bf h}\in{\mathbb{C}}^{M\times 1}$ denote the narrowband channel vector associated with $M$ densely deployed antennas, and $Q$ denote the number of transmit pilots within a coherence-time frame. The received signal ${y}_{q}\in{\mathbb{C}}$ at the BS in timeslot $q$ is modeled as
@@
where ${\bf w}_{q}\in{\mathbb{C}}^{M\times 1}$ is the observation vector at the BS, $s_{q}$ the pilot transmitted by the user, and ${\bf z}_{q}\sim\mathcal{C}\mathcal{N}\!\left({\bf 0}_{M},\sigma^{2}{\bf I}_{M}\right)$ the additive white Gaussian noise. As the observation vector ${\bf w}_{q}$ affects both the desired signal ${\bf h}s_{q}$ and the noise ${\bf z}_{q}$, its power has no impact on the estimation accuracy. Therefore, without any loss of generality, we can normalize ${\bf w}_{q}$ to $\|{\bf w}_{q}\|^{2}_{2}=1$, which reshapes the noise distribution to ${\bf w}_{q}^{H}{\bf z}_{q}\sim\mathcal{C}\mathcal{N}\!\left(0,\sigma^{2}\right)$. For ease of notation, we assume that $s_{q}=1$ for all $q\in\{1,\cdots,Q\}$.
As shown in Fig. 2, in each timeslot, the BS reconfigures its observation vector to obtain sufficient observations. Considering the total $Q$ timeslots for pilot transmission, we get
@@
where ${\bf y}=\left[{y}_{1},\cdots,{y}_{Q}\right]^{\rm T}$, ${\bf{W}}=\left[{\bf{w}}_{1},\cdots,{\bf{w}}_{Q}\right]$ stands for the observation matrix, and ${\bf z}:=\left[{\bf w}^{H}_{1}{\bf z}_{1},\cdots,{\bf w}^{H}_{Q}{\bf z}_{Q}\right]^{\rm T}\sim\mathcal{CN}(0,\sigma^{2}{\mathbf{I}}_{Q})$. The goal of channel estimation is to recover the channel $\bf h$ from the received pilot $\bf y$. In particular, the Bayesian estimation criterion promises to minimize the MSE, $\mathsf{E}\|{\mathbf{h}}-\hat{{\mathbf{h}}}\|_{2}^{2}$, by designing the reconstructor $\hat{{\mathbf{h}}}$ as a function of $\bf y$ and our prior knowledge to ${\mathbf{h}}$. In existing literature, a significant amount of effort has endeavored to design advanced reconstructors for making the estimation of ${\mathbf{h}}$ as accurate as possible, such as CS algorithms and deep-learning based methods [27, 28, 29, 30, 31, 33, 34, 35]. However, the design of the observation matrix, ${\mathbf{W}}$, for channel estimation is largely unexplored. It is worth noting that the extremely high spatial correlation exhibited by DAS channels leads to a remarkable deviation of the channel covariance matrix from the identity matrix. This deviation highlights the significance of aligning the observation matrix with the channel’s subspace for boosting the estimation accuracy. Motivated by this fact, this paper proposes to optimize the reconstructor and the observation matrix jointly, i.e.,
$\min_{\hat{{\mathbf{h}}},{\mathbf{W}}\in\mathcal{W}}\mathsf{E}\|{\mathbf{h}}-\hat{{\mathbf{h}}}\|_{2}^{2},$ (3)
where $\mathcal{W}$ stands for the feasible set of the observation matrix ${\mathbf{W}}$. For instance, let’s consider the $q$-th observation vector, ${\mathbf{w}}_{q}$, from the matrix ${\mathbf{W}}$. In the case of amplitude-and-phase controllable scenarios, the constraint imposed on ${\mathbf{w}}_{q}$ is $\|{\mathbf{w}}_{q}\|_{2}^{2}=1$; while in the case of phase-only controllable scenarios [44], each entry of ${\mathbf{w}}_{q}$ is subject to the constant modulus constraint, i.e., $|w_{q,m}|=\frac{1}{\sqrt{M}}$ for $m\in\{1,2,\cdots,M\}$.
Figure 2: The frame structure for DAS channel estimation.

### v2

# DAS_CE_Supp

@@ Abstract
By deploying a large number of antennas with sub-half-wavelength spacing in a compact space, dense array systems (DASs) can fully unleash the multiplexing and diversity gains of limited apertures. To acquire these gains, accurate channel state information acquisition is necessary but challenging due to the large antenna numbers. To overcome this obstacle, this paper reveals that designing the observation matrix to exploit the high spatial correlation of DAS channels is crucial for realizing optimal/near-optimal channel estimation. Specifically, we prove that the observation matrix design for channel estimation is equivalent to a time-domain duality of point-to-point multiple-input multiple-output precoding, except for the change in the total power constraint on the precoding matrix to the pilot-wise discrete power constraint on the observation matrix. A novel ice-filling algorithm is proposed to design amplitude-and-phase controllable observation matrices, and a majorization-minimization algorithm is proposed to address the phase-only controllable case. Particularly, we prove that the ice-filling algorithm can be interpreted as a “quantized” water-filling algorithm, wherein the latter’s continuous power-allocation process is converted into the former’s discrete pilot-assignment process. To support the sub-optimality of the proposed designs, we provide comprehensive analyses on the achievable mean square errors and their asymptotic expressions. Finally, numerical results confirm that our proposed designs achieve the near-optimal channel estimation performance and outperform existing approaches significantly.
@@ I Introduction
From 3G to 5G, antenna array systems play an irreplaceable role in wireless communications [1, 2]. By strategically configuring multiple antennas, an antenna array can constructively manipulate the radiated signals, allowing for enhanced wireless coverage, improved spectral efficiency, and reduced power consumption [3, 4, 5]. The performance of arrays continuously improves with the number of antennas [6, 7]. However, in practical scenarios, the space available for antenna deployment is usually limited, which restricts the performance gains endowed by antenna arrays [8]. To achieve a better performance with a spatially-limited aperture, dense array systems have attracted extensive attentions in recent years [9].
@@
Generally, DASs represent a series of array technologies that massive sub-wavelength antennas are densely arranged within a compact space. Unlike the conventional arrays with half-wavelength antenna spacing $\lambda/2$, the antenna spacing of DASs is much smaller, such as $\lambda/6$ [10], $\lambda/8$ [11], $\lambda/10$ [12], or even $\lambda/23$ [13]. With an increased number of sub-channels, DASs promise to achieve high array gains and fully exploit the multiplexing and diversity gains of limited apertures [9]. Besides, DASs can also reduce the grating lobes and provide high performance for large values of oblique angles of incidence [14]. Some works also report their advantages of super-bandwidth and super-directivity [15, 16, 17]. The typical DAS realizations include holographic multiple-input multiple-output (MIMO) [18], reconfigurable intelligent surfaces [19], fluid antenna systems [20, 21], graphene-based nano-antenna arrays [22, 23], radio-frequency lens [24], and so on. For example, in [10], massive sub-wavelength patches of $\lambda/6$-spacing are densely printed on a holographic MIMO surface to generate multiple beams flexibly. In [25], meta-elements sized of $\lambda/2.5\times\lambda/17.5$ are seamlessly integrated onto a feeding microstrip to achieve reconfigurable holographic surface (RHS) beam steering. In [13], an RIS composed of antennas sized of $\lambda/23\times\lambda/23$ is designed and fabricated for terahertz communications.
@@
The performance gains of antenna arrays are realized via the constructive beamformers enabled by their phase-shifters (PSs) and radio frequency (RF) chains. To realize effective beamforming and signal demodulation, the acquisition of accurate channel state information (CSI) is essential, particularly in situations where the number of RF chains is smaller than the number of massive antennas. Up to now, many estimators have been proposed to acquire the CSI of large antenna arrays. Their fundamental principle involves receiving pilot signals using observation matrices and recovering the channel by advanced estimation algorithms. For example, when the available pilot length is larger than the antenna number, the classical non-parametric algorithms are widely used, including least square (LS) estimator and minimum mean square error (MMSE) estimator. Leveraging the property of channel sparsity, compressed sensing (CS)-based channel estimators are widely studied to improve the estimation accuracy and reduce the pilot overhead [26, 27], such as the orthogonal matching pursuit (OMP)-based estimator [28, 29], the message passing (MP)-based estimator [30, 31, 32], and the gridless sparse signal reconstructor [33]. By training neural networks with a large amounts of channel data, the deep learning approaches are also utilized to realize data-driven channel estimators [34, 35, 36]. Besides, beam alignment techniques [37, 38, 39], including beam sweeping and hierarchical beam training, have also been widely explored to acquire the implicit CSI with low pilot overhead.
@@
Although most of existing channel estimators can be adopted in DASs, they fail to fully exploit the high spatial correlation of DAS channels, thereby leaving a remarkable performance gap from the optimal estimator. Specifically, this spatial correlation is attributed to the fact that, the extremely-dense deployment of DAS antennas significantly increases the similarity of radio waves impinging on antenna ports and aggravates the mutual-coupling effect between adjacent antenna circuits. From the mathematical perspective, this fact implies that the covariance matrices of DAS channels are no longer diagonal, but highly structured and under-determined. For diagonal covariance matrices, a uniform or random pilot-sensing strategy is sufficient to achieve the optimality of estimation [40, 41]. Thereby, in conventional channel estimations, the observation matrices for receiving the pilot signals are either generated randomly or set as predefined codebooks, such as the Discrete Fourier Transform (DFT) matrix and the identity matrix [26, 27, 28, 33, 30, 29, 31, 34, 42, 37, 38, 32, 36, 39]. In contrast, since the correlation matrix of DAS channels are full of structural features, it is believed that their observation matrices can be tightly aligned with these features in pilot transmission. This structured pilot-sensing process has a high potential to remarkably boost the channel estimation accuracy in DASs [43, 44, 45].
@@
To achieve this goal, our work represents the first attempt on designing and analyzing the optimal/near-optimal observation matrices in the DAS channel estimation. Our main contributions are summarized as follows.
@@
  • Mutual-information-maximization (MIM) based framework: From the perspective of MIM, we propose to design the DAS’s observation matrix by minimizing the information uncertainty between the received pilots and wireless channels. The formulated MIM problem for observation matrix design is shown to be a time-domain duality of point-to-point MIMO precoding design. In particular, the multi-RF-chain resources for MIMO precoding is transformed to the multi-timeslot-pilot resources for channel estimation, and the total power constraint on the precoding matrix is switched to the pilot-wise discrete power constraint on the observation matrix. Motivated by this finding, we show that the ideal (but unachievable in practice) observation matrix could be constructed by the water-filling principle, which involves the use of the eigenvectors of channel covariance matrix as well as optimal power allocation.
@@
  • Majorization minimization (MM) enabled observation matrix design: We then apply our framework for DAS channel estimation to the situation when only the phases of the observation matrix’s weights are controllable while their amplitudes are fixed, making both the water-filling and ice-filling algorithms invalid. To address this challenge, we propose a MM algorithm for designing the observation matrix. Its novelty lies in replacing the primal non-convex MIM problem with a series of tractable approximate subproblems having analytical solutions, which are solved in an alternating optimization manner. Numerical simulations demonstrate that its channel estimation performance is comparable to the ice-filling algorithm.
  • Performance analysis: Comprehensive analyses on the achievable mean square errors (MSEs) are provided to validate the effectiveness of the proposed designs. We prove that, compared to the random observation matrix design, the ice-filling algorithm can significantly improve the estimation accuracy by the order of the ratio between the number of antennas and the rank of channel covariance. Furthermore, based on the derived quantization nature of ice-filling, the MSE gap between water-filling and ice-filling is proved to decay quadratically with the pilot length, demonstrating the near-optimality of the ice-filling algorithm. In addition, the close-form MSEs under imperfect channel covariance are also derived. The superiority of the proposed algorithms is further validated using extensive numerical results.
The remainder of the paper is organized as follows. The system model is presented in Section II. The proposed MIM-based framework for observation matrix design is given in Section III. The ice-filling and MM algorithms are elaborated in Section IV. Then, the MSEs are analyzed in Section V. Simulation results are carried out in Section VI, and finally the conclusions are drawn in Section VII.
@@ II System Model
Figure 1: An illustration of uplink channel estimation for an DAS.
As illustrated in Fig. 1, we consider the narrowband channel estimation of an uplink single-input multiple-output (SIMO) system. The base station (BS) employs an $M$-antenna DAS, which comprises an analog combining structure supported by one RF chain, to receive the pilots from a single-antenna user.
Let ${\bf h}\in{\mathbb{C}}^{M\times 1}$ be the channel vector and $Q$ the number of transmit pilots within a coherence-time frame. The received signal ${y}_{q}\in{\mathbb{C}}$ at the BS in timeslot $q$ is modeled as
@@
where ${\bf w}_{q}\in{\mathbb{C}}^{M\times 1}$ is the observation vector at the BS, $s_{q}$ the pilot transmitted by the user, and ${\bf z}_{q}\sim\mathcal{C}\mathcal{N}\!\left({\bf 0}_{M},\sigma^{2}{\bf I}_{M}\right)$ the additive white Gaussian noise. As the observation vector ${\bf w}_{q}$ affects both the desired signal ${\bf h}s_{q}$ and the noise ${\bf z}_{q}$, its power has no impact on the estimation accuracy. Therefore, without any loss of generality, we can normalize ${\bf w}_{q}$ to $\|{\bf w}_{q}\|^{2}_{2}=1$, which reshapes the noise distribution to ${\bf w}_{q}^{H}{\bf z}_{q}\sim\mathcal{C}\mathcal{N}\!\left(0,\sigma^{2}\right)$. Two practical implementations of the observation vector ${\bf w}_{q}$ will be discussed in our research: the amplitude-and-phase controllable combiner and the phase-only controllable combiner. The former connects each antenna to the RF chain using one PS and one low noise amplifier (LNA), enabling adjustment of both the amplitude and phase of the coefficients of ${\bf w}_{q}$. The latter deploys a single PS between each antenna and the RF chain, allowing only phase adjustment of the coefficients of ${\bf w}_{q}$ and enforcing a constant modulus constraint on the observation vector, i.e., $|w_{q,m}|=\frac{1}{\sqrt{M}}$ for $m\in\{1,2,\cdots,M\}$ [26].
Considering the total $Q$-timeslot pilot transmission, the received signal vector ${\bf y}:=\left[{y}_{1},\cdots,{y}_{Q}\right]^{\rm T}$ is expressed as
@@
where $s_{q}$ is assumed to be 1 for all $q\in\{1,\cdots,Q\}$, ${\bf{W}}=\left[{\bf{w}}_{1},\cdots,{\bf{w}}_{Q}\right]$ stands for the observation matrix, and ${\bf z}:=\left[{\bf w}^{H}_{1}{\bf z}_{1},\cdots,{\bf w}^{H}_{Q}{\bf z}_{Q}\right]^{\rm T}$. As the power of ${\mathbf{w}}_{q}$ is normalized to 1, we have ${\bf z}\sim\mathcal{CN}({\mathbf{0}}_{Q},\sigma^{2}{\mathbf{I}}_{Q})$. The objective of channel estimation is to recover the channel $\bf h$ from the received pilot $\bf y$. As a DAS possesses highly correlated channels across different antenna ports, we adopt Bayesian regression to recover ${\mathbf{h}}$. Suppose the channel is generated from a Gaussian process $\mathcal{CN}({\mathbf{0}}_{M},\bm{\Sigma}_{\bf h})$, where the covariance matrix $\bm{\Sigma}_{\bf h}\in\mathbb{C}^{M\times M}$, also called prior kernel, characterizes our prior knowledge to the channel correlation. The joint probability distribution of $[{\mathbf{h}}^{T},{\mathbf{y}}^{T}]^{T}$ satisfies
[…]
Thereby, the estimated channel, $\hat{{\mathbf{h}}}$, is determined by the posterior mean of ${\mathbf{h}}$ given observation ${\mathbf{y}}$:
${{\bm{\mu}}}_{{\mathbf{h}}|{\mathbf{y}}}={{\bf{\Sigma}}_{\bf{h}}}{{\bf{W}}}{\left({\bf W}^{H}{\bf{\Sigma}}_{\bf{h}}{\bf W}+\sigma^{2}{\bf I}_{Q}\right)^{-1}}{{\bf{y}}}.$ (4)
The posterior covariance matrix of ${\mathbf{h}}$, also called posterior kernel, is accordingly derived as
Notably, the posterier kernel, characterizing the channel estimation error, is a function of the prior kernel ${{\bf{\Sigma}}_{\bf{h}}}$ and the observation matrix ${\mathbf{W}}$. Due to the extremely high spatial correlation exhibited by DAS channels, the prior kernel ${{\bf{\Sigma}}_{\bf{h}}}$ would remarkably deviate from the identity matrix. This deviation indicates that aligning the observation matrix ${\mathbf{W}}$ with the channel’s subspace could be greatly helpful in reducing the estimation error. Motivated by this fact, this paper concentrates on the design of observation matrix ${\mathbf{W}}$, by considering both the amplitude-and-phase controllable and phase-only controllable analog combiners, to boost the channel estimation performance.